Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
4x + 3y = -45.5
-5x + 8y = -43

Answer by mananth(16946)

4x+3y= -45.5 .............1
-5x+8y=-43 .............2
Eliminate y
multiply (1)by -8
Multiply (2) by 3
-32x-24 y=364
-15x+24 y=-129
Add the two equations
-47 x = 235
/-47
x=-5.00
plug value of x in (1)
4x+3y=-45.5
-20+3y=-45.5
3y=-45.5 +20
3y=-25.5
y=-8.50
(-5,-8.50)